The sum of two numbers is 31. Their difference is 9. What are the two numbers?
Answer by checkley71(8403)

x+y=31 or x=31-y
x-y=9
now replace x by (31-y) & solve for y
31-y-y=9
31-2y=9
-2y=9-31
-2y=-22
y=-22/-2
y=11 answer.
x+11=31
x=31-11
x=20 answer.
proof.
20+11=31